OpenStudy (anonymous):

Can Someone PLEASE explain to me rational and irrational numbers? I DO NOT IN ANY WAY GET THIS!! I have looked everywhere :( Please explain! Thanks in advance.

OpenStudy (amistre64):

what is your idea of a ratioed number?

OpenStudy (tkhunny):

Rational -- Think RATIO

OpenStudy (anonymous):

@amistre64 Honestly, I have no clue. I'm completely clueless here.

OpenStudy (phi):

The Ancient Greeks did not get it either

OpenStudy (amistre64):

if we have an integer ratio of n:k then the ratioed number is n/k any number that looks like a fraction where n and k are integers, k not 0 is a rational number

OpenStudy (amistre64):

if a number cannot be written, as the ratio of 2 integers, then it is not a rational number

OpenStudy (amistre64):

you had to be working in base 60 to get it :)

OpenStudy (anonymous):

Okay... that makes a bit of sense... I mean, I know more than I did XD Thanks! @amistre64

OpenStudy (amistre64):

youre welcome. the more you work with the number sets the easier itll become.

OpenStudy (tkhunny):

Ratio 1/2 Rational 3/5 Rational 23/19 Rational If is can be expressed as a ratio of two integers, it's Rational. What's an Integer?

OpenStudy (ghostgate):

A rational number is any number that can be written as a fraction, on the other hand, an irrational number is a number that can't be written as a fraction. For instance: 4.75 = 19/4 = Rational 0.5 = 1/2 = Rational 1.125 = 9/8 = Rational Whereas: Pi = Irrational sqrt2 = Irrational etc. I hope this helps in someway! Have a great day! ;) {---Ghostgate---}